OpenStudy (anonymous):

Multiply the polynomials (x^3+x-2)4x

OpenStudy (anonymous):

4x^4 +4x^2 - 8x (:

OpenStudy (anonymous):

Use FOIL. Remember when you multiply exponents, you add them. ex) x^2 (x^2) = x^ (2+2) = x^4 x^3 (x^5) = x ^ (3+5)= x^8 This would be the same as 4x(x^3)+4x(x)-4x(2) = 4x^4+4x^2-8x